11/23/99 Owl shift summary

Ashkan_Alavi-Harati (alavi@fnal.gov)
Tue, 23 Nov 1999 07:02:54 -0600

Shift Summary for Owl shift, 23/Nov/99

On Shift: Ashkan, M. Barrio [MB: 0-3:30; AA: 3:30-7]

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Scheduled down time for RunII studies

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:
none.

========================================================================
Accesses during shift:
none.

========================================================================
Overview of shift:
Cosmic data taken.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15034 Ongoing cosmic run from previous shift

We have had some SDAQ and pipeline problems, all
probably related to the fact that we are using a
fake spill signal.

pipeline: tried to reset it, but upon launching of
pipe.tcl (on atev2 as user da_ktev3) I get a
warning winda saying that running the script as
da_ktev3 could cause serious trouble, so I just
leave it alone to it's own devices.

SDAQ alarm: probably a fake spill signal
problem... Ashkan should revise the status when
he gets in.

2:28 Got a coupla ktevana errors. Had to run kfrend
twice before I could silence Elliot. But then
again, I'm fortunate that Val hasn't "panicked"
so far.

3:30 Ashkan Replaces Miguel for the rest of the shift.
SDAQ process which require info from the EPICURE
and/or spill signal are down. The fake spill
signal is not doing SDAQ any good. In this
situation one can ignore errors from 'sad' and
'ktscaler' process. Fresh start wouldn't help.

3:55 A trigger error, KFREND resolves.

6:15 SDAQ is getting the correct spill number all
of a sudden! All process up and running.

6:18 Recovering from TeV runII studies, preparing for
shut down. No beam to fixed target till this
evening.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift